OpenStudy (anonymous):

Quadratic equations that can be solved using the quadratic formula can also be solved by factoring.

OpenStudy (anonymous):

true or false?

OpenStudy (anonymous):

yes...........true.......but it will be diffcult as sometimes answers might be in irrational numbers, imaginary unit, decimals, .....

OpenStudy (anonymous):

true

OpenStudy (anonymous):

oops its actually sometimes always never

OpenStudy (anonymous):

always

hero (hero):

sometimes

OpenStudy (anonymous):

its called the box method

OpenStudy (anonymous):

in what case, we cant solve HERO

hero (hero):

I think joe math might be able to explain it better than I can

OpenStudy (anonymous):

but, even equations with roots imaginary unit, fractions and decimals can be solved by factorization........difference is it will be time taking and harder

OpenStudy (anonymous):

The answer the teacher wants is sometimes, but the actual answer is always. put sometimes though. If this is high school, most high school teachers dont consider things like:\[(x-\frac{1+\sqrt{5}}{2})(x-\frac{1-\sqrt{5}}{2})=x^2-x-1\] to be proper.

hero (hero):

I was going to post a response similar to that

OpenStudy (anonymous):

But the actual answer is a polynomial can always be factored in the complex numbers.

OpenStudy (anonymous):

so, answer is always, isnt it

OpenStudy (anonymous):

you can also use a graphing calculator by going into y= putting in quadratic equation pressing 2nd then trace(will come up as calc) go to zero answer three questions (left,right,guess) then u get your answer

hero (hero):

I know that all polynomials are factorable, like joe said, it's the answer that the teacher wants that counts

OpenStudy (anonymous):

yeah, im my experience with high school teachers ( in the US anyways), the answer is sometimes.

hero (hero):

This is where edit button is needed. I meant to say all quadratics*

OpenStudy (anonymous):

in*

hero (hero):

I'm going to bed. It's getting very late. Bye everyone